by weight, one alloy is 75%copper and 25%zinc. Another alloy is 42%copper and 58%zinc. How many grams of each or these are required to make 200 grams of an alloy that is 60% copper and 40% zinc?
Answer by Paul(988)

0.75x+0.42(200-x)=0.6(200)
0.75x+84-0.42x=120
0.33x=36
x=109
200-109=91
Hence for 75$ copper and 25% zine it's 109 grams and for other it's 91grams